OpenStudy (anonymous):

Four friends played cards. Their final scores were –6, –3, 0, and 1. What is the mean absolute deviation of their scores?

hartnn (hartnn):

whats the mean ?

OpenStudy (anonymous):

2.5

OpenStudy (anonymous):

-2 is the mean i think

hartnn (hartnn):

correct :) you know what to do next?

OpenStudy (anonymous):

No lol

hartnn (hartnn):

take the difference of each score with the mean! so -6 - (-2) = .. ? -3 - (-2) = ..? and so on...

OpenStudy (anonymous):

-4 and -1???

hartnn (hartnn):

yes, but we use absolute values, so 4 and 1. 0 and 1 are still left 0-(-2) = 1-(-2) =.. ?

OpenStudy (anonymous):

2 and 3

hartnn (hartnn):

now take the average of these 1,4,2,3

OpenStudy (anonymous):

10 and i divide it right? (If yes its 2.5)

hartnn (hartnn):

good! 2.5 is correct :)

OpenStudy (anonymous):

Thx :)

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

I got a.......

OpenStudy (anonymous):

80%

OpenStudy (anonymous):

I passed it! YAY!

hartnn (hartnn):

CONGRATULATIONS!! :D

OpenStudy (anonymous):



hartnn (hartnn):

\[ \begin{array}l\color{red}{\text{w}}\color{orange}{\text{e}}\color{#E6E600}{\text{l}}\color{green}{\text{c}}\color{blue}{\text{o}}\color{purple}{\text{m}}\color{purple}{\text{e}}\color{red}{\text{ }}\color{orange}{\text{^}}\color{#E6E600}{\text{_}}\color{green}{\text{^}}\color{blue}{\text{}}\end{array} \]